OpenStudy (anonymous):

Two mechanics worked on a car. The first mechanic worked for 5 hours, and the second mechanic worked for 15 hours. Together they charged a total of $1275 . What was the rate charged per hour by each mechanic if the sum of the two rates was $155 per hour?

OpenStudy (anonymous):

Help Please.

OpenStudy (anonymous):

@optiquest Help Please

OpenStudy (anonymous):

5*m1 + 15*m2 = 1275 m1+m2 =155 m1,m2 are rates for each mechanic

OpenStudy (anonymous):

@optiquest I dont understand

OpenStudy (anonymous):

The first mechanic worked for 5 hours mechanic 1 rate times 5 hours second mechanic worked for 15 hours. mechanic 2 times 15 hours Together they charged a total of $1275 mechanic 1 rate times 5 hours + mechanic 2 times 15 hours = $1275

OpenStudy (anonymous):

What was the rate charged per hour by each mechanic if the sum of the two rates was $155 per hour? mechanic 1 rate + mechanic 2 rate = $155 this is a system of two equations which you need to solve

OpenStudy (wolf1728):

5f + 15s = 1,275 f + s = 155 You have to solve for 2 unknowns so you need 2 equations.

OpenStudy (anonymous):

@wolf1728 so I solve for s?

OpenStudy (anonymous):

is it 275? for the first and

OpenStudy (wolf1728):

You need to find f and s Do you know how to solve these equations? No it is NOT 275 for the first

OpenStudy (anonymous):

@wolf1728 No, I don't

OpenStudy (wolf1728):

You need to eliminate 1 of the unknowns first. A) 5f + 15s = 1,275 B) f + s = 155 Multiply Equation B by -5 and so we get A) 5f + 15s = 1,275 B) -5f -5s = -775 Next we add both equations giving 10s = 500 s = 50

OpenStudy (anonymous):

@wolf1728 after this then what do I do?

OpenStudy (wolf1728):

Now that the value of s has been found, you next solve for f.

OpenStudy (anonymous):

so plug in 50 for both

OpenStudy (anonymous):

@FlashBT3

OpenStudy (wolf1728):

no - f + s = 155 so if s=50 then f= what?

OpenStudy (anonymous):

55

OpenStudy (wolf1728):

f +50 = 155 So, f = ?

OpenStudy (anonymous):

105

OpenStudy (wolf1728):

Yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Listen to @wolf1728

OpenStudy (anonymous):

he know what he is talking about.

OpenStudy (anonymous):

So first mechanic is 50 and second is 105

OpenStudy (wolf1728):

thanks flash

OpenStudy (anonymous):

@wolf1728

OpenStudy (anonymous):

yes you are correct